OpenStudy (anonymous):

A triangle has a perimeter of 50. If 2 of its sides are equal and the third side is 5 more than the equal sides, what is the length of the third side?

OpenStudy (anonymous):

let a be one side, b second side, c is third side. since two sides are equal a=b. third side is 5 more than the equal sides so c = 5+a. the sum should be 50 a + a + (5+a) = 50 3a +5 = 50 3a = 50-5 3a = 45 a=15 now you can solve for the third side

OpenStudy (anonymous):

2x + (x + 5) = 50 3x = 45 Now solve for x....

OpenStudy (anonymous):

thank you..

OpenStudy (anonymous):

...a lot

OpenStudy (anonymous):

Welcome dear..